# Learning Network Dismantling without Handcrafted Inputs

### v1

@@ Appendix C Generating Synthetic Networks
Random Network Models
[…]
LPA model starts with a clique of $m+1$ nodes. Then, at each step, a new node forms connections to $m$ existing nodes with the probability of choosing an existing node to form a connection being proportional to $d_{i}+m(\gamma-3)$, where $d_{i}$ is the degree of an existing node $v_{i}$. This results in a scale-free network with a power-law degree distribution $P(d)\sim d^{-\gamma}$ and an average degree of $\langle d\rangle=2m$. When generating using the Copying Model, we also start with a clique of $m+1$ nodes. Each of the $m$ edges of a newly added node is formed by either i) connecting to a randomly selected existing node, with probability $\alpha=\frac{2-\gamma}{1-\gamma}$, or ii) connecting to one neighbor of a randomly selected existing node, with probability $1-\alpha$. This also results in a scale-free network with a power-law degree distribution $P(d)\sim d^{-\gamma}$ and an average degree of $\langle d\rangle=2m$. (LPA and Copying Model generate fundamentally different structures given the same parameters, due to their different mechanisms of edge formation.) When generating using the ER model, for all pairs of nodes, an edge will be formed with probability:
$p=\frac{[2N-(m+1)]m}{N(N-1)},$
leading to a network whose degree distribution is Binomial (well-approximated by Poisson) with an average degree of $\langle d\rangle=m$.
Degree-Preserving Rewiring
@@
[…]
Appendix D MIND Training
Training of MIND is conducted on an Ubuntu 22.04 server with an Intel(R) Xeon(R) w9-3475X CPU (512 GB RAM) and an RTX 5000 Ada GPU (32 GB VRAM). The same set of hyperparameters (listed in Table 3), mostly aligned with the hyperparameters for discrete SAC in (Huang et al. 2022), is used across all experiments. The Adam optimizer is used for updating the neural networks. To estimate the state-action value, we apply a forgetting factor of $\lambda=0.99$, resulting in the following:
which slightly discounts future AUC contributions but effectively prevents unbounded growth of the cumulative AUC for large networks. To stabilize the bootstrapping estimation of $Q(G_{t},v_{i})$, we use target Q-networks that remain fixed for 8000 steps before being updated. By setting the target Q-network update factor to 1, the target networks are synchronized with the main Q-networks every 8000 steps. The pseudo-code for the training process of MIND is provided in Algorithm 1.
Parameter Value
Total number of steps $s_{\text{total}}$ $8\times 10^{6}$
Replay buffer size $|\mathcal{D}|$ $2\times 10^{6}$
Learning rate for Q-network $3\times 10^{-4}$
Learning rate for policy network $3\times 10^{-4}$
Batch size for updating $|\mathcal{B}|$ 512512512
Start learning $s_{\text{start}}$ 100000100000100000
Target Q-network update factor 111
Forgetting factor $\gamma$ 0.990.990.99
Policy network update frequency 444
Target network update frequency 800080008000
Node embedding size $F$ 4
Number of heads $H$ 444
Message-passing iterations $K$ 666
$\mathrm{MLP}_{\sigma}$ and $\mathrm{MLP}_{\nu}$ structure 16,[32],116,[32],116 , [ 32 ] , 1
$\mathrm{MLP}_{\theta}$ and $\mathrm{MLP}_{\phi}$ $\text{structure}^{*}$ 196,[256,256],1196,[256,256],1196 , [ 256 , 256 ] , 1
Table 3: The hyperparameters of MIND. For Neural Networks, the first number is the input size, the numbers in [][$\cdot$][ ⋅ ] are the size of the hidden layers, and the last number is the output size.
Appendix E Statistics on the Real Networks
In Table 4, we present key statistics of the real-world networks used in the Experiments section. Specifically, we report the total number of nodes $|V|$, average degree, degree-assortativity coefficient (within the range [1,1][-1,1][ - 1 , 1 ]), and modularity (within the range [1,1][-1,1][ - 1 , 1 ]) of each network. The table shows that these networks span a wide range of sizes (from 128128128 to 1.41.41.4 million nodes) and average degrees (from 222 to 333333). Many real networks are disassortative (or exhibit negative degree-assortativity), whereas certain networks—mostly technological networks—have very high modularity.
Appendix F Per-Network Results on Real-World Networks
In Table 5, we report numerical values of the relative AUC of the dismantling curve, given by every method on all of the real-world networks. These values are used to generate the scatter and bar plots in Fig 2.
@@ Appendix F Per-Network Results on Real-World Networks
Table 4: Statistics on the real networks.
@@
Table 5: Performance of different methods on real networks.

### v2

@@ Appendix C Generating Synthetic Networks
Random Network Models
To synthesize each training network, we randomly select one of three generation models—LPA, Copying Model, or ER—with equal probability. The network size $|V|$ is sampled uniformly from the range of $100$ to $200$. The average degree is determined by the parameter $m$ drawn randomly either from the set $\{1,6,8,10\}$ with probability $1/3$, or from the set $\{2,3,4,5\}$ with probability $2/3$. For Copying Model and LPA networks, a degree distribution (power-law) exponent $\gamma\in[2,4]$ is also selected uniformly at random.
The LPA model starts with a clique of $m+1$ nodes. Then, at each step, a new node forms connections to $m$ existing nodes with the probability of choosing an existing node to form a connection being proportional to $d_{i}+m(\gamma-3)$, where $d_{i}$ is the degree of an existing node $v_{i}$. This results in a scale-free network with a power-law degree distribution $P(d)\sim d^{-\gamma}$ and an average degree of $\langle d\rangle=2m$. When generating using the Copying Model, we also start with a clique of $m+1$ nodes. Each of the $m$ edges of a newly added node is formed by either i) connecting to a randomly selected existing node, with probability $\alpha=\frac{2-\gamma}{1-\gamma}$, or ii) connecting to one neighbor of a randomly selected existing node, with probability $1-\alpha$. This also results in a scale-free network with a power-law degree distribution $P(d)\sim d^{-\gamma}$ and an average degree of $\langle d\rangle=2m$. (LPA and Copying Model generate fundamentally different structures given the same parameters, due to their different mechanisms of edge formation.) When generating using the ER model, for all pairs of nodes, an edge will be formed with probability $p=\frac{m}{n-1}$, leading to a network whose degree distribution is Binomial (well-approximated by Poisson) with an average degree of $\langle d\rangle=m$.
Degree-Preserving Rewiring
@@
With a coin flip, we choose to label the nodes either i) randomly, or ii) in the order of their degree (nodes with the same degree are given consecutive integers as labels). The former choice will lead to the process of controlling the modularity of the network, and the latter choice will lead to the control of the network’s degree-assortativity. This controlling of the modularity/degree-assortativity is achieved by the iterative application of degree-preserving edge rewirings. The target (label-)assortativity coefficient is then sampled uniformly from the set $\{0.05,0.15,0.2,0.25,0.3,0.4,0.5\}$ and a random (negative or positive) sign is applied to the coefficient. Starting with any network, we perform a randomly picked label-assortative (resp. label-disassortative rewiring) if the network’s current label-assortativity is below (resp. above) the target coefficient. Double-edges rewirings, in the appropriate direction, are iteratively performed until the network’s assortativity (computed corresponding to the chosen node labeling) matches the target value within tolerance or the maximum number of rewiring attempts is reached. Note that the resulting rewired networks will always be connected.
@@
Appendix D MIND Training
MIND is trained on an Ubuntu 22.04 server with an Intel(R) Xeon(R) w9-3475X CPU (512 GB RAM) and an RTX 5000 Ada GPU (32 GB VRAM). On this machine, the complete training process took approximately 90 h, which is comparable to the training time reported for the RL-based baseline method FINDER. MIND requires substantially less training time than the other machine learning-based baseline, GDM, which relies on brute-force dismantling of training networks with factorial complexity in the number of nodes. The pseudo-code of the training process of MIND is presented in Algorithm 1.
The same set of hyperparameters (listed in Table 3), mostly aligned with the hyperparameters for discrete SAC in (Huang et al. 2022), is used across all experiments. The Adam optimizer is used for updating the neural networks. To estimate the state-action value, we apply a forgetting factor of $\lambda=0.99$, resulting in the following:
which slightly discounts future AUC contributions but effectively prevents unbounded growth of the cumulative AUC for large networks. To stabilize the bootstrapping estimation of $Q(G_{t},v_{i})$, we update the target Q-networks intermittently (every 8000 steps). By setting the target Q-network update factor to 1, in the update event, the target networks are synchronized with the main Q-networks. The number of message-passing iterations is set to $K=6$. Although increasing the number of message-passing iterations $K$ enlarges the receptive field of the GNN encoder in MIND, it leads to higher computational complexity and increased training difficulty. Also, most real-world networks exhibit small-world properties with small diameters, motivating our choice of a small $K$. MIND uses a relatively lightweight model with 120 k trainable parameters. This is approximately 100 times smaller than the best-performing baseline, GDM, with 13 M parameters. The final embedding dimension of MIND is 96 (after concatenation of MIND-MP), compared to 1500 in GDM.
Parameter Value
Total number of steps $s_{\text{total}}$ $8\times 10^{6}$
Replay buffer size $|\mathcal{D}|$ $2\times 10^{6}$
Learning rate for Q-network $3\times 10^{-4}$
Learning rate for policy network $3\times 10^{-4}$
Batch size for updating $|\mathcal{B}|$ $512$
Start learning $s_{\text{start}}$ $100000$
Target Q-network update factor 11
Forgetting factor $\gamma$ $0.99$
Policy network update frequency 44
Target network update frequency $8000$
Node embedding size $F$ 4
Number of heads $H$ 44
Message-passing iterations $K$ 66
$\mathrm{MLP}_{\sigma}$ and $\mathrm{MLP}_{\nu}$ structure $16,[32],1$
$\mathrm{MLP}_{\theta}$ and $\mathrm{MLP}_{\phi}$ $\text{structure}^{*}$ $196,[256,256],1$
Table 3: The hyperparameters of MIND. For Neural Networks, the first number is the input size, the numbers in $\cdot$] are the size of the hidden layers, and the last number is the output size.
Appendix E Statistics on the Real Networks
In Table 4, we present key statistics of the real-world networks used in the Experiments section. Specifically, we report the total number of nodes $|V|$, average degree, degree-assortativity coefficient (within the range [1,1][-1,1]), and modularity (within the range [1,1][-1,1]) of each network. The table shows that these networks span a wide range of sizes (from $128$ to $1.4$ million nodes) and average degrees (from $2$ to $33$). Many real networks are disassortative (or exhibit negative degree-assortativity), whereas certain networks—mostly technological networks—have very high modularity.
Since the message-passing operations in MIND are confined to local neighborhoods, its computational complexity scales linearly with the number of nodes, while allowing parallelization on modern GPUs. Such efficiency enables MIND to scale to huge real-world networks. For example, MIND requires only 6.8 hours to dismantle the largest real network, Hyves, which contains 1.5 million nodes, compared to 9.8 hours for GDM and 12.2 hours for FINDER.
Appendix F Detailed Dismantling Results on Real-World Networks
In Table 5, we report numerical values of the relative AUC of the dismantling curve, given by every method on all of the real-world networks. The results shown in Fig 2 are generated from the data reported in Table 5. MIND’s better performance over the state-of-the-art baselines is statistically significant, with $p$-values of <0.008<0.008 for the Wilcoxon test comparing MIND versus each baseline.
Appendix G Limitation and Future Works
Despite the state-of-the-art performance, dismantling networks with huge diameters is the weakness of MIND. This is a common weakness of all common message-passing GNNs arising from their limited message-passing reach, and remains to be solved while respecting the linear computational complexity. Another potential future work is the RL training framework. In this manuscript, we did not delve into the effects of the RL model on the dismantling performance. We based our choice of the RL algorithm on observing the performance of a number of recent methods in elementary tests. Due to its off-policy learning, SAC uses the observed network dismantling steps more efficiently in training. Exploiting the most recent advances in RL and/or tailoring RL models (e.g., leveraging domain knowledge to reduce the policy search space) for the purpose of network dismantling is a potential direction for future work. Last but not least, the application of the ideas presented in this manuscript can be investigated to derive GNN-based solutions to general complex network problems.
@@ Appendix G Limitation and Future Works
Table 4: Statistics on the real-world networks.
@@
Table 5: Detailed dismantling results on real-world networks.